# Empirical and computer-aided robustness analysis of long-step and accelerated methods in smooth convex optimization

### v1

@@ 1.1 Inexact gradient methods
$\|d_{k}-\nabla f(x_{k})\|\leq\delta\|\nabla f(x_{k})\|\text{, with }\delta\in[0,1)\text{ the relative size of inexactness.}$.
@@ 1.4 Introduction to the PEP methodology
Solving the optimization problem (PEP) gives the tight worst-case performance of a first-order algorithm $\mathcal{A}$ on the class of functions $\mathcal{F}$, such as the class of smooth convex functions. This worst-case performance is said to be tight, as it is exactly equal to the worst-case rate. The last constraint bounds the difference in objective function value between the initial point $x_{0}$ and the optimal solution $x_{\ast}$, as it is well known that in most situations, the performance of a first-order method cannot be sensibly assessed without such a constraint. The distance between $x_{0}$ and $x_{\ast}$ can also be considered, leading to potentially different rates, but in this work, we focus on the objective function difference, leading to the convergence rate $\tau_{N}$ introduced earlier. With the numerical resolution of the PEP, reformulated as a tractable semidefinite program, this methodology provides insights into worst-case behaviors. It enables the derivation of tight performance guarantees, which are often difficult to obtain analytically.
@@ 2 Worst-case Results
The above-described PEP methodology allows us to get the numerical worst-case performance of a given method with fixed parameters. Given that the schedule of step sizes presented in Section 1.3 is already fixed, fixing different values of inexactness $\delta$ allows us to get the theoretical convergence rate $\tau_{N}$.
@@ 3 Empirical Results
We apply this logistic regression model to a binary image classification task, using a dataset of handwritten images of digits 0 and 1 [7]. (We also ran similar tests on another dataset involving phishing websites [1], and observed similar conclusions.)
Each image is represented as a vector of pixel values $x_{k}$, and the associated label $y_{k}\in\{0,1\}$ indicates whether the image shows a zero ($y_{k}=0$) or a one ($y_{k}=1$). We denote by $K$ the total number of training images.
@@ 4 Conclusion
Future work The encouraging results from this work open up the possibility of future work. The use of compression of the mantissa could be explored on more complex problems, such as small neural networks with sigmoid activation functions, generalizing the concept of logistic classification. Furthermore, a stochastic analysis of the compressed gradient descent, instead of only considering the worst-case compression error, would also be interesting. Indeed, on average, errors caused by compression are much smaller than the worst-possible ones.

### v2

@@ 1.1 Inexact gradient methods
$\|d_{k}-\nabla f(x_{k})\|\leq\delta\|\nabla f(x_{k})\|\text{, with relative inexactness level }\delta\in[0,1)$ (2)
@@ 1.4 Introduction to the PEP methodology
Solving the optimization problem (PEP) gives the tight worst-case performance of a first-order algorithm $\mathcal{A}$ on the class of functions $\mathcal{F}$, such as the class of smooth convex functions. This worst-case performance is said to be tight, as it is exactly equal to the worst-case rate. The last constraint bounds the difference in objective function value between the initial point $x_{0}$ and the optimal solution $x_{\ast}$, as it is well known that in most situations, the performance of a first-order method cannot be sensibly assessed without such a constraint. The distance between $x_{0}$ and $x_{\ast}$ can also be considered, leading to potentially different rates. Still, in this work, we focus on the objective function difference, leading to the convergence rate $\tau_{N}$ introduced earlier.
To solve problem (PEP), which contains an infinite-dimensional variable (the function $f$), it must first be reformulated using interpolation conditions characterizing the class of functions $\mathcal{F}$, as described in [13]. Inexactness of the gradient can also be dealt with in this setting by writing condition (2) as an additional constraint [5, 12]. This leads to a tractable semidefinite program to compute the tight convergence rate $\tau_{N}$ and obtain insights about worst-case behaviors.
@@ 2 Worst-case Results
The above-described PEP methodology allows us to get the numerical worst-case performance of a given method with fixed parameters. In practice we used the PEPit toolbox [8] with the MOSEK solver [3]. Given that the schedule of step sizes presented in Section 1.3 is already fixed, fixing different values of inexactness $\delta$ allows us to get the theoretical convergence rate $\tau_{N}$.
@@ 3 Empirical Results
We apply this logistic regression model to a binary image classification task, using a dataset of handwritten images of digits 0 and 1 [9]. (We also ran similar tests on another dataset involving phishing websites [1], and observed similar conclusions.) Each image is represented as a vector of pixel values $x_{k}$, and the associated label $y_{k}\in\{0,1\}$ indicates whether the image shows a zero ($y_{k}=0$) or a one ($y_{k}=1$). We denote by $K$ the total number of training images.
@@ 4 Conclusion
Future work Encouraging results from this work open up some possibility of future work. The use of compression of the mantissa could be explored on more complex problems, such as small neural networks with sigmoid activation functions, generalizing the concept of logistic classification. Furthermore, a stochastic analysis of compressed gradient descent, instead of only considering the worst-case compression error, would also be interesting. Indeed, on average, errors caused by compression are much smaller than the worst-possible ones.